# X-GS: An Extensible Open Framework Unifying 3DGS Architectures with Downstream Multimodal Models

### 10 Mar 2026 v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, subsequently extending into numerous spatial AI applications. However, most existing 3DGS methods are isolated, focusing on specific domains such as online SLAM, semantic enrichment, or 3DGS for unposed images. In this paper, we introduce X-GS, an extensible open framework that unifies a broad range of techniques to enable real-time 3DGS-based online SLAM enriched with semantics, bridging the gap to downstream multimodal models. At the core of X-GS is a highly efficient pipeline called X-GS-Perceiver, capable of taking unposed RGB (or optionally RGB-D) video streams as input to co-optimize geometry and poses, and distill high-dimensional semantic features from vision foundation models into the 3D Gaussians. We achieve real-time performance through a novel online Vector Quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a highly parallelized pipeline design. The semantic 3D Gaussians can then be utilized by vision-language models within the X-GS-Thinker component, enabling downstream tasks such as object detection, zero-shot caption generation, and potentially embodied tasks. Experimental results on real-world datasets showcase the efficacy, efficiency, and newly unlocked multimodal capabilities of the X-GS framework.
X-GS: An Extensible Open Framework Unifying 3DGS Architectures with Downstream Multimodal Models
Yueen Ma1, Irwin King1 Department of Computer Science and Engineering, The Chinese University of Hong Kong1 {yema21, king}@cse.cuhk.edu.hk
@@ 1 Introduction
Figure 1: X-GS is an extensible open framework that unifies previously isolated domains, including pose-free 3DGS, 3DGS-based SLAM, Semantic GS, and VLMs for 3DGS. X-GS-Perceiver achieves real-time 3DGS-based online SLAM with semantics using an online VQ module, GPU-accelerated grid-sampling, and highly parallelized scheduling. X-GS-Thinker bridges the resulting semantic 3DGS representation with downstream multimodal models to execute complex, language-driven tasks such as open-vocabulary 3D object detection, caption generation, and potentially embodied tasks.
@@
Despite this rapid progress, these advancements typically operate in isolation. To combine their respective strengths, we propose the first work to unify these disparate 3DGS techniques from a comprehensive framework perspective: the X-GS framework. Standing for an eXtensible open framework for 3DGS, the “X” serves as a placeholder prefix to denote distinct 3DGS paradigms, such as Pose-Free-GS, SLAM-GS, Semantic-GS, and VLM-for-GS. This suggests that future developments in these different directions can be assimilated by X-GS for continued improvement, directly showcasing its extensibility. Built upon this unified foundation, X-GS achieves real-time 3DGS-based online SLAM with rich semantic enrichment while simultaneously bridging the gap to downstream VLMs.
Table 1: Comparison of X-GS with representative 3DGS methods.
Method RGB-only RGB-D Pose-free Online-SLAM Real-time Semantics VLM-Integration
@@ 2 Related Work
Figure 2: Overview of the X-GS framework. X-GS-Perceiver synergizes a memory-efficient Vector Quantization (VQ) module, grid-based semantic supervision, and an asynchronous parallelized pipeline to perform SLAM and distill semantics simultaneously in an online fashion, operating in real time at $\sim$15 FPS. As an open framework, it accommodates both RGB-only and RGB-D inputs, and can flexibly integrate various Vision Foundation Models (VFMs). Furthermore, the X-GS-Thinker component is extensible to different multimodal models, enabling a wide range of downstream tasks.
@@ 3 Method
We propose X-GS, an extensible open framework designed to unify previously isolated 3DGS advancements. The architecture is divided into two primary subsystems. First, the perception module, X-GS-Perceiver, ingests unposed monocular or RGB-D video streams to continuously co-optimize the 3D Gaussian map and camera poses, while distilling high-dimensional foundation model features into the Gaussians as semantic channels. To ensure real-time performance, it utilizes online vector quantization, grid-sampling, and parallel scheduling. Second, different multimodal models can serve as the X-GS-Thinker. By utilizing the enriched 3DGS representation produced by X-GS-Perceiver, the Thinker unlocks advanced capabilities such as text-prompted 3D object detection, zero-shot scene captioning, and embodied tasks.
@@ 3.1 Preliminaries
$\mathbf{T}_{t}^{*}=\underset{\mathbf{T}_{t}}{\arg\min}\ \mathcal{L}_{\text{track}}(I_{t},D_{t}\mid\mathbf{T}_{t},\Theta)$ (2)
@@
$\Theta^{*}=\underset{\Theta}{\arg\min}\sum_{k\in\mathcal{K}}\mathcal{L}_{\text{map}}(I_{k},D_{k}\mid\Theta,\mathbf{T}_{k})$ (3)
@@ 3.2 X-GS-Perceiver
We develop a set of modules that add semantic enrichment capabilities to the online geometric and photometric optimization pipeline. Naively attaching semantic features to 3D Gaussians has proven to be heavily memory- and compute-intensive. To ensure our pipeline still fulfills the real-time requirement, we devise three key speedup techniques: an online vector quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a meticulously designed pipeline that fully leverages highly parallelized scheduling.
@@ 4.1 Implementation Details
4.2 Contrastive VLM Results
To evaluate the visual and semantic fidelity of our system, we present qualitative comparisons of scene reconstruction and open-vocabulary object localization in Figure 3. First, we observe that the proposed X-GS-Perceiver module maintains exceptional geometry and appearance tracking; the rendered RGB reconstructions exhibit high visual fidelity and closely match the ground truth (GT) images. Second, we evaluate the distilled semantic field. Despite running in an online fashion at real-time speeds, our framework successfully distills semantic information from the VFMs of SAM and CLIP. The integration of the VQ module and grid-sampled supervision effectively balances speed and quality. Finally, we demonstrate the practical utility of our semantically enriched 3D scenes. By querying the distilled semantic field using text prompts, our framework computes per-Gaussian similarity scores to accurately localize specific objects in 3D space. This confirms that X-GS operates not merely as an efficient SLAM system, but successfully constructs a deeply comprehensible, queryable 3D semantic field ready for complex downstream tasks.
4.3 Generative VLM Results
@@
Figure 4: Qualitative results of X-GS for 3D scene caption generation.
@@ 4.4 Computational Footprint Analysis
Table 2: Computational footprint analysis of X-GS-Perceiver. We employ SigLIP and evaluate the pipeline on a single NVIDIA V100 GPU.
@@ 5 Conclusion
In this paper, we introduced X-GS, an extensible framework that unifies isolated 3DGS domains—including pose-free tracking, online SLAM, and semantic 3DGS—into a single cohesive system. To achieve real-time performance, we proposed the X-GS-Perceiver, which overcomes computational bottlenecks using an online VQ module with EMA updates, GPU-accelerated grid sampling, and highly parallelized scheduling. Furthermore, we demonstrated the framework’s capability for multimodal reasoning through the X-GS-Thinker, a module that bridges our semantic 3D representations with downstream VLMs and VLAs. By natively supporting complex tasks such as open-vocabulary 3D object detection, scene captioning, and embodied AI planning, X-GS establishes a robust, modular foundation for real-time spatial computing.
@@
While X-GS establishes a highly extensible framework, its multimodal reasoning capabilities require more comprehensive evaluation across diverse embodied and 3D vision-language benchmarks. Currently, the interface with the X-GS-Thinker operates in a modular fashion; future work will explore end-to-end fine-tuning of the downstream VLMs natively within the 3DGS representations. Additionally, as demonstrated by recent systems like FreeSplat Wang et al. (2024b) and SplatTalk Thai et al. (2025), there is a growing trend toward feed-forward 3DGS generation and reasoning. Integrating such feed-forward mechanisms into X-GS could significantly reduce the current optimization overhead. Furthermore, leveraging emerging 3D vision foundation models—such as DUSt3R Wang et al. (2024a), MASt3R Leroy et al. (2024), or VGGT Wang et al. (2025)—could provide stronger geometric priors for mapping, further enhancing the robustness of our pose-free tracking pipeline. Finally, accommodating dynamic scenes using techniques such as 4DGS represents a highly promising direction for real-world embodied AI applications.

### 12 Mar 2026 v2

# X-GS: An Extensible Open Framework for Perceiving and Thinking via 3D Gaussian Splatting

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, subsequently extending into numerous spatial AI applications. However, most existing 3DGS methods operate in isolation, focusing on specific domains such as pose-free 3DGS, online SLAM, and semantic enrichment. In this paper, we introduce X-GS, an extensible open framework consisting of two major components: the X-GS-Perceiver, which unifies a broad range of 3DGS techniques to enable real-time online SLAM and distill semantic features; and the X-GS-Thinker, which interfaces with downstream multimodal models. In our implementation of the Perceiver, we integrate various 3DGS methods through three novel mechanisms: an online Vector Quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a highly parallelized pipeline design. The Thinker accommodates vision-language models and utilizes the resulting 3D semantic Gaussians, enabling downstream applications such as object detection, caption generation, and potentially embodied tasks. Experimental results on real-world datasets demonstrate the efficiency and newly unlocked multimodal capabilities of the X-GS framework.
X-GS: An Extensible Open Framework for Perceiving and Thinking via 3D Gaussian Splatting
Yueen Ma13, Zenglin Xu23, Irwin King1 The Chinese University of Hong Kong1, Fudan University2, Shanghai Academy of AI for Science3 {yema21, king}@cse.cuhk.edu.hk, zenglinxu@fudan.edu.cn
{NoHyper}This is a work in progress.
@@ 1 Introduction
Figure 1: X-GS is an extensible open framework that unifies previously isolated domains in 3D-GS. X-GS-Perceiver achieves real-time 3DGS-based online SLAM with semantics using an online VQ module, GPU-accelerated grid-sampling, and highly parallelized scheduling. X-GS-Thinker bridges the resulting 3D semantic Gaussians with diverse downstream multimodal models and tasks.
@@
Despite this rapid progress, these advancements typically operate in isolation. To combine their respective strengths, we propose the first work to unify these disparate 3DGS techniques from a comprehensive framework perspective: the X-GS framework, as illustrated in Figure 1. Standing for an eXtensible open framework for 3DGS, the “X” serves as a placeholder prefix denoting distinct 3DGS paradigms, such as RGBD-3D-GS, pose-free-GS, online-GS, SLAM-GS, semantic-GS, and VLM-for-GS. This design ensures that future developments in these different directions can be assimilated into X-GS for continued improvement, directly showcasing its extensibility. Built upon this unified foundation, X-GS achieves real-time 3DGS-based online SLAM with semantic distillation while simultaneously bridging the gap to downstream VLMs.
Table 1: Comparison of X-GS with representative 3DGS methods. Gen-VLM: integration with generative VLMs.
Method RGB-only RGB-D Pose-free Online-SLAM Real-time Semantics Gen-VLM
@@ 2 Related Work
Figure 2: Overview of the X-GS framework. X-GS-Perceiver synergizes a memory-efficient Vector Quantization (VQ) module, grid-based semantic supervision, and an asynchronous parallelized pipeline to perform SLAM and distill semantics simultaneously in an online fashion, executing in real time at $\sim$15 FPS. As an open framework, it accommodates both RGB-only and RGB-D inputs, and can flexibly integrate various Vision Foundation Models (VFMs). Furthermore, the X-GS-Thinker component is extensible to different multimodal models, enabling a wide range of downstream tasks.
@@ 3 Method
We propose X-GS, an extensible open framework designed to unify previously isolated 3DGS advancements. As illustrated in Figure 2, the framework is divided into two primary components. First, the perception component, X-GS-Perceiver, ingests unposed monocular or RGB-D video streams to continuously co-optimize the 3D Gaussians and camera poses, while simultaneously distilling high-dimensional foundation model embeddings into the Gaussians as semantic features. To ensure real-time performance, it utilizes online vector quantization, grid sampling, and parallel scheduling. Second, the X-GS-Thinker component accommodates various multimodal models. By leveraging the 3D semantic Gaussians produced by the Perceiver, the Thinker unlocks advanced capabilities such as text-prompted 3D object detection, zero-shot scene captioning, and potentially embodied tasks.
@@ 3.1 Preliminaries
$C(x\mid\mathbf{T},\Theta)=\sum_{i=1}^{N}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j}),$ (1)
@@
$\mathbf{T}_{t}^{*}=\underset{\mathbf{T}_{t}}{\arg\min}\ \mathcal{L}_{\text{track}}(I_{t},D_{t}\mid\mathbf{T}_{t},\Theta).$ (2)
@@
$\Theta^{*}=\underset{\Theta}{\arg\min}\sum_{k\in\mathcal{K}}\mathcal{L}_{\text{map}}(I_{k},D_{k}\mid\Theta,\mathbf{T}_{k}).$ (3)
@@ 3.2 X-GS-Perceiver
We develop a set of modules that add semantic enrichment capabilities to the online geometric and photometric optimization pipeline. Naively attaching semantic features to 3D Gaussians has proven to be heavily memory- and compute-intensive. To ensure our pipeline still fulfills the real-time requirement, we devise three key speedup techniques: an online Vector Quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a meticulously designed pipeline that fully leverages highly parallelized scheduling. Specifically, the VQ module reduces the feature dimensionality of each pixel within the rasterized semantic map, while the grid-sampling scheme reduces its resolution (height and width). Finally, parallelized scheduling optimizes the overall execution at the module level.
@@ 4.1 Implementation Details
4.2 X-GS-Perceiver Results
To evaluate the visual and semantic fidelity of our system, we present qualitative comparisons of scene reconstruction and open-vocabulary object localization in Figure 3. First, we observe that the proposed X-GS-Perceiver module maintains exceptional geometry and appearance tracking; the rendered RGB reconstructions exhibit high visual fidelity and closely match the ground truth (GT) images. Second, we evaluate the distilled semantic field. Despite running in an online fashion at real-time speeds, our framework successfully distills semantic information from the VFMs of SAM and CLIP. The integration of the VQ module and grid-sampled supervision effectively balances speed and quality. Finally, we demonstrate the practical utility of our semantically enriched 3D scenes. By querying the distilled semantic field using text prompts, our framework computes per-Gaussian similarity scores to accurately localize specific objects in 3D space. This confirms that X-GS operates not merely as an efficient SLAM system, but successfully constructs a deeply comprehensible, queryable 3D semantic field ready for complex downstream tasks. We also include results for when X-GS receives additional depth map inputs in Figure 4. A quantitative comparison of these settings is provided in Table 2.
Figure 4: Qualitative results of X-GS using RGB-D inputs. Depth map is optional and not the required input by X-GS-Perceiver.
Table 2: Comparison of RGB-only and RGB-D settings on TUM fr3/office. Rendering metrics are reported after color refinement.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE RMSE $\downarrow$
RGB-only 27.039 0.8651 0.1698 0.01368
RGB-D 27.304 0.8797 0.1280 0.00952
4.3 X-GS-Thinker Results
@@
Figure 5: Qualitative results of X-GS for 3D scene caption generation. The VLM generates captions in a zero-shot setting by taking 3D semantic Gaussians as input. Note that only a portion of the generated content is shown.
@@ 4.4 Computational Footprint Analysis
Table 3: Computational footprint analysis of X-GS-Perceiver. We employ SigLIP as its VFM and evaluate the pipeline on a single NVIDIA V100 GPU.
@@ 5 Conclusion
In this paper, we introduced X-GS, an extensible framework that unifies previously isolated 3DGS domains—including pose-free GS, online SLAM, and semantic GS—into a single cohesive system. To achieve real-time performance for X-GS-Perceiver, we overcome computation bottlenecks using an online VQ module with EMA updates, GPU-accelerated grid sampling, and highly parallelized scheduling. Furthermore, we demonstrated the framework’s capability for multimodal reasoning through X-GS-Thinker, a module that bridges our 3D semantic Gaussians with downstream VLMs. By natively supporting multimodal tasks such as open-vocabulary object detection, scene captioning, and embodied tasks, X-GS establishes a robust, modular foundation for spatial reasoning.
@@
While X-GS establishes a highly extensible framework, its multimodal reasoning capabilities require more comprehensive evaluation across diverse embodied and 3D vision-language benchmarks. First, the interface with the X-GS-Thinker currently operates in a modular fashion; future work will explore end-to-end fine-tuning of the downstream VLMs natively within the 3DGS representations Thai et al. (2025). Second, as demonstrated by recent systems like FreeSplat Wang et al. (2024b), SplatTalk Thai et al. (2025), EA3D Zhou et al. (2025), and EmbodiedSplat Lee et al. (2026), there is a growing trend toward feed-forward architectures for 3DGS. Integrating such feed-forward mechanisms into X-GS could significantly reduce the current optimization overhead, but such models also require significant training cost for different VFMs. Third, leveraging emerging 3D vision foundation models—such as DUSt3R Wang et al. (2024a), MASt3R Leroy et al. (2024), or VGGT Wang et al. (2025)—could provide stronger geometric priors for mapping, further enhancing the robustness of our pose-free tracking pipeline. Finally, accommodating dynamic scenes using techniques such as 4DGS Wu et al. (2024); Gao et al. (2025) represents a highly promising direction for real-world embodied AI applications.